# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ Abstract
We study a new class of MDPs that employs multinomial logit (MNL) function approximation to ensure valid probability distributions over the state space. Despite its benefits, introducing the non-linear function raises significant challenges in both computational and statistical efficiency. The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes. While this result attains the same rate in $K$ as linear cases, the method requires storing all historical data and suffers from an $\mathcal{O}(K)$ computation cost per episode. Moreover, the quantity $\kappa$ can be exponentially small in the worst case, leading to a significant gap for the regret compared to linear function approximation. In this work, we first address the computational and storage issue by proposing an algorithm that achieves the same regret with only $\mathcal{O}(1)$ cost. Then, we design an enhanced algorithm that leverages local information to enhance statistical efficiency. It not only maintains an $\mathcal{O}(1)$ computation and storage cost per episode but also achieves an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$, nearly closing the gap with linear function approximation. Finally, we establish the first lower bound for MNL function approximation, justifying the optimality of our results in $d$ and $K$.
@@ 1 Introduction
While these studies make important steps toward understanding RL with function approximation, there are still some challenges to be solved. In linear function approximation, transitions are assumed to be linear in specified feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to state $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, a transition function is a probability distribution over states, i.e., the magnitude falls within the range of $[0,1]$, and the aggregation equals $1$. For certain feature mappings, linear transitions may not yield a valid probability distribution with arbitrary $\theta$ and $\mu$ as show by (Hwang and Oh, 2023). While there are also some works exploring generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they made assumptions over value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous works and ours in terms of the regret, computation cost and storage cost. Here $\kappa$ and $\kappa^{*}$ are exponentially small problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost and storage cost per episode indicate the dependence on episode count $k$.
Reference Regret Com. Sto. MDP model
@@
UCRL-MNL-OL (Theorem 1) $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
UCRL-MNL-LL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
@@
Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ / / infinite action space
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit function approximation has been proposed by Hwang and Oh (2023) recently. Despite its benefits, the introduction of non-linear functions raises significant challenges in both computational and statistical efficiency. Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret. While matching the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ as linear function approximation in terms of $K$ (Zhou et al., 2021; He et al., 2023), their method requires storing all historical data and the computational cost per episode grows linearly with the episode count, which is expensive. Moreover, the quantity $\kappa$ can be exponentially small, leading to a significant gap for the regret compared to the linear cases. Thus, a natural question arises:
Is MNL function approximation more difficult than linear function approximation for RL?
In this paper, we answer this question by making significant advancements in both computational and statistical efficiency for MDPs with MNL function approximation, nearly closing the gap with linear function approximation. Table 1 presents a comparison between previous studies and our work. Specifically, our contributions are summarized as follows:
  • We first propose the UCRL-MNL-OL algorithm based on a variant of Online Newton Step, attaining an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret with $\mathcal{O}(1)$ computational and storage costs per episode. This result matches the best-known regret by Hwang and Oh (2023), yet achieves the same computational and storage efficiency as linear function approximation (Jin et al., 2020; Zhou et al., 2021).
  • As the quantity $\kappa$ can be exponentially small, we propose an enhanced algorithm UCRL-MNL-LL, which leverages the local information to improve the statistical efficiency. It not only maintain $\mathcal{O}(1)$ computational and storage costs but achieve an improved regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+d^{2}H^{2}\kappa^{-1})$. The higher-order term matches the optimal regret $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ for linear mixture MDPs (Zhou et al., 2021) and linear MDPs (He et al., 2023) in $d$ and $K$, differing only by an $\mathcal{O}(H^{1/2})$ factor.
@@
  • We establish the first lower bound for MDPs with MNL function approximation by introducing a reduction to the logistic bandit problem. We prove a lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for MDPs with infinite action space. Though this does not constitute a strict lower bound for the finite actions case studied in this work, it suggests that our result may be optimal with respect to $d$ and $K$.  [^1] After the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) improved the lower bound to $\Omega(dH^{3/2}\sqrt{K})$. This confirms that our result is indeed optimal with respect to $d$ and $K$.
From a technical perspective, inspired by recent advances in logistic bandits (Zhang et al., 2016; Faury et al., 2020; Zhang and Sugiyama, 2023), we observe that the negative log-likelihood function is exponentially concave, which motivates us to apply online Newton step to estimate the unknown transition parameter in an online manner. Moreover, we employ the Bernstein-like inequalities and the self-concordant-like property (Bach, 2010) of the log-loss to achieve the better dependence on $\kappa$.
@@
Organization.  We introduce the related work in Section 2 and present the setup in Section 3. Then, we design an algorithm with constant computational and storage costs in Section 4. Next, we present an algorithm with improved regret guarantee in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space limits, we defer all proofs to the appendixes.
Notations.  Denote by $[n]$ the set $\{1,\ldots,n\}$ and use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$. For $X,Y\in\mathbb{R}^{d\times d}$, $X\succeq Y$ means $X-Y$ is positive semi-definite. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. The $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors.
@@ 2 Related Work
RL with Generalized Linear Function Approximation.  There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach which models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is characterized using a multinomial logit model. We significantly improve upon their results by providing computationally and statistically more efficient algorithms.
@@
RL with General Function Approximation.  There have also been some works studies RL with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions to ensure valid probability distributions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
@@
Multinomial Logit Bandits.  There are two types of multinomial logit bandits studied in the literature: the single-parameter model, where the parameter is a vector (Cheung and Simchi-Levi, 2017) and multiple-parameter model, where the parameter is a matrix (Amani and Thrampoulidis, 2021). We focus on the single-parameter model, which are more relevant to our setting. The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{K})$, where $K$ denotes the number of rounds in bandits. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ in the uniform revenue setting. Most prior methods required storing all historical data and faced computational challenge. To address this issue, the most recent work by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs building on recent advances in multiple-parameter model (Zhang and Sugiyama, 2023). Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa K}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{K}+\kappa^{-1})$ under uniform and non-uniform rewards respectively. However, although the underlying models of MNL bandits and MDPs share similarities, the challenges they present differ substantially, and techniques developed for MNL bandits cannot be directly applied to MNL MDPs. For example, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward. Thus, it is necessary to design new algorithms tailored for MDPs to address these unique challenges.
@@ 3 Problem Setup
Inhomogeneous, Episodic MDPs.  An inhomogeneous, episodic MDP instance can be denoted by a tuple $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
@@
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$. The Bellman equation for policy $\pi$ and Bellman optimality equation are given by
$Q_{h}^{\pi}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{\pi}\right](s,a),\quad Q_{h}^{*}(s,a)=r_{h}(s,a)+\left[\mathbb{P}_{h}V_{k,h+1}^{*}\right](s,a).$
Learning Protocol.  In the online MDP setting, the learner interacts with the environment without the knowledge of the transition $\{\mathbb{P}_{h}\}_{h=1}^{H}$. As learning rewards is no harder than transitions, we assume the reward $\{r_{h}\}_{h=1}^{H}$ is known. The interaction proceeds in $K$ episodes. At the beginning of episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as
@@
Multinomial Logit (MNL) Mixture MDPs.  Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model (including bilinear and low-rank MDPs) can induce a valid probability distribution over the state space, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. We introduce the definition of MNL mixture MDPs below.
@@
Remark 1.
There are many cases that even when the state space is very large, the maximum number of reachable states $U$ is small. This phenomenon is common in situations where the next state is close to the current state. An illustrative example is the RiverSwim problem (Strehl and Littman, 2004).
@@
An MDP instance $M=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
@@
While the work of Hwang and Oh (2023) focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$), we address the more general inhomogeneous setting, allowing transitions to vary across different stages.
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. We introduce the key problem-dependent quantities for this problem firstly introduced by Hwang and Oh (2023).

### v3

@@ Abstract
We study a new class of MDPs that employs multinomial logit (MNL) function approximation to ensure valid probability distributions over the state space. Despite its significant benefits, incorporating the non-linear function raises substantial challenges in both statistical and computational efficiency. The best-known result of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret upper bound, where $\kappa$ is a problem-dependent quantity, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes. However, we observe that $\kappa^{-1}$ exhibits polynomial dependence on the number of reachable states, which can be as large as the state space size in the worst case and thus undermines the motivation for function approximation. Additionally, their method requires storing all historical data and the time complexity scales linearly with the episode count, which is computationally expensive. In this work, we propose a statistically efficient algorithm that achieves a regret of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$, eliminating the dependence on $\kappa^{-1}$ in the dominant term for the first time. We then address the computational challenges by introducing an enhanced algorithm that achieves the same regret guarantee but with only constant cost. Finally, we establish the first lower bound for this problem, justifying the optimality of our results in $d$ and $K$.
@@ 1 Introduction
While these studies make important steps toward understanding RL with function approximation, there are still challenges to be solved. In linear function approximation, transitions are assumed to be linear in feature mappings, such as $\mathbb{P}(s^{\prime}|s,a)=\phi(s^{\prime}|s,a)^{\top}\theta^{*}$ for linear mixture MDPs and $\mathbb{P}(s^{\prime}|s,a)=\phi(s,a)^{\top}\mu^{*}(s^{\prime})$ for linear MDPs. Here $\mathbb{P}(s^{\prime}|s,a)$ is the probability from state $s$ to $s^{\prime}$ taking action $a$, $\phi(s^{\prime}|s,a)$ and $\phi(s,a)$ are feature mappings, $\theta^{*}$ and $\mu^{*}(s^{\prime})$ are unknown parameters. However, the transition function is a probability distribution over states, meaning its values must lie within $[0,1]$ and sum to 1. Thus, the linearity assumption is restrictive and hard to satisfy in practice. An algorithm designed for linear MDPs could break down entirely if the underlying MDP is not linear (Jin et al., 2020). While some works explore generalized linear (Wang et al., 2021) and general function approximation (Russo and Roy, 2013; Foster et al., 2021; Chen et al., 2023), they focus on function approximation for value functions rather than transitions, hence do not tackle this challenge.
Table 1: Comparison between previous works and ours in terms of the regret and computational cost (including storage and time complexity) . Here $\kappa$ and $\kappa^{*}$ are problem-dependent quantities defined in Assumption 1, $d$ is the feature dimension, $H$ is the episode length and $K$ is the number of episodes. The computational cost per episode only highlight the dependence on the episode count $k$.
Reference Regret Storage Time MDP model
@@
UCRL-MNL-LL (Theorem 1) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(k)$ $\mathcal{O}(k)$ inhomogeneous
UCRL-MNL-OL (Theorem 2) $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$ $\mathcal{O}(1)$ $\mathcal{O}(1)$ inhomogeneous
@@
Lower Bound  (Corollary 1) $\Omega(dH\sqrt{K\kappa^{*}})$ infinite action space
Towards addressing the limitation of linear function approximation, a new class of MDPs that utilizes multinomial logit function approximation has been proposed by Hwang and Oh (2023) recently. Such formula also aligns better with models like neural networks (LeCun et al., 2015), which inherently respect the probabilistic constraints through a softmax layer and allow for greater expressive power. However, though it offers promising benefits, the introduction of non-linear functions introduces significant challenges in both statistical and computational efficiency. Specifically, the best-known approach of Hwang and Oh (2023) has achieved an $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ regret, where $\kappa$ is a problem-dependent quantity that measures the effective non-linearity over the entire parameter space, $d$ is the feature dimension, $H$ is the episode length, and $K$ is the number of episodes. Unfortunately, as we show in Claim 1, it holds that $\kappa^{-1}>U^{2}$, where $U$ denotes the maximum number of reachable states, which equals to the size of the state space $S$ in the worst case. This undermines the core motivation for function approximation, which aims to mitigate dependence on large state and action spaces. Furthermore, the method requires storing all historical data, and its time complexity per episode grows linearly with the episode count (i.e., $\mathcal{O}(k)$ at episode $k$). Thus, a natural question arises:
Is it possible to design both statistically and computationally efficient algorithms
for RL with MNL function approximation?
In this work, we answer this question affirmatively for the class of MNL mixture MDPs where the transition is parameterized by a multinomial logit function. Our contributions are listed as follows:
  • For statistical efficiency, we propose the UCRL-MNL-LL algorithm, which attains a regret bound of $\widetilde{\mathcal{O}}(dH^{2}\sqrt{K}+\kappa^{-1}d^{2}H^{2})$. Our result significantly improves upon the $\widetilde{\mathcal{O}}(\kappa^{-1}dH^{2}\sqrt{K})$ rate of Hwang and Oh (2023), making the first time to achieve a $\kappa$-independent dominant term (note that the lower-order term still scales with $\kappa^{-1}$, but does not depend on $K$, making it acceptable). To achieve this, we propose a tighter confidence set based on a new Bernstein-type concentration (Périvier and Goyal, 2022) instead of the standard Hoeffding-type concentration, and exploit the self-concordant-like property (Bach, 2010) of the log-loss function to better use local information.
  • For computational efficiency, we propose the UCRL-MNL-OL algorithm, which enjoys the same regret bound as UCRL-MNL-LL, but with only constant storage and time complexity per episode. This is enabled by recognizing that the negative log-likelihood function is exponentially concave, which motivates the use of online mirror descent with a specifically tailored local norm (Zhang and Sugiyama, 2023) to replace the standard maximum likelihood estimation. Furthermore, we construct the optimistic value function by incorporating a closed-form bonus term through a second-order Taylor expansion, thus avoiding the need to solve a non-convex optimization problem.
@@
  • We establish the first lower bound for MNL mixture MDPs by introducing a reduction to the logistic bandit problem. We prove a problem-dependent lower bound of $\Omega(dH\sqrt{K\kappa^{*}})$ for infinite action setting, where $\kappa^{*}$ is an another problem-dependent quantity that measures the effective non-linearity over around the ground truth parameter. Though this does not constitute a strict lower bound for the finite action case studied in this work, it suggests that our result may be optimal in $d$ and $K$.  [^1] After the submission of our work to arXiv (Li et al., 2024a), a follow up work by Park et al. (2024) proved a lower bound of $\Omega(dH^{3/2}\sqrt{K})$ for the finite action setting. This confirms that our result is optimal in $d$ and $K$.
Table 1 provides a comparison between our work and previous studies, focusing on regret and computational costs, including both storage and time complexity.
@@
Organization.  We introduce the related work in Section 2 and present the setup in Section 3. Then, we design a statistically efficient algorithm in Section 4. Next, we present an algorithm that achieves both statistical and computational efficiency in Section 5. Finally, we establish the lower bound in Section 6. Section 7 concludes the paper. Due to space limits, we defer all proofs to the appendixes.
Notations.  We use $[x]_{[a,b]}$ to denote $\min(\max(x,a),b)$. For a vector $\mathbf{x}\in\mathbb{R}^{d}$ and positive semi-definite matrix $A\in\mathbb{R}^{d\times d}$, denote $\lVert\mathbf{x}\rVert_{A}=\sqrt{\mathbf{x}^{\top}A\mathbf{x}}$. For a strictly convex and continuously differentiable function $\psi:\mathcal{W}\mapsto\mathbb{R}$, the Bregman divergence is defined as $\mathcal{D}_{\psi}\left(\mathbf{w}_{1},\mathbf{w}_{2}\right)=\psi\left(\mathbf{w}_{1}\right)-\psi\left(\mathbf{w}_{2}\right)-$ $\left\langle\nabla\psi\left(\mathbf{w}_{2}\right),\mathbf{w}_{1}-\mathbf{w}_{2}\right\rangle$. We use the notation $\mathcal{O}(\cdot)$ to indicate different types of dependencies depending on the context. For regret analysis, $\mathcal{O}(\cdot)$ omits only constant factors. For computational costs, we use $\mathcal{O}(\cdot)$ to solely highlight the dependence on the number of episode as this is the primary factor influencing the complexity. Additionally, we employ $\widetilde{\mathcal{O}}(\cdot)$ to hide all polylogarithmic factors.
@@ 2 Related Work
RL with Generalized Linear Function Approximation.  There are recent efforts devoted to investigating function approximation beyond the linear models. Wang et al. (2021) investigated RL with generalized linear function approximation. Notably, unlike our approach that models transitions using a generalized linear model, they apply this approximation directly to the value function. Another line of works (Chowdhury et al., 2021; Li et al., 2022; Ouhamma et al., 2023) has studied RL with exponential function approximation and also aimed to ensure that transitions constitute valid probability distributions. The MDP model can be viewed as an extension of bilinear MDPs in their work while our setting extends linear mixture MDPs. These studies are complementary to ours and not directly comparable. Moreover, these works also enter the computational and statistical challenges arising from non-linear function approximation that remain to be addressed. The most relevant work to ours is the recent work by Hwang and Oh (2023), which firstly explored a similar setting to ours, where the transition is characterized using a multinomial logit model. We significantly improve upon their results by providing statistically and computationally more efficient algorithms.
@@
RL with General Function Approximation.  There have also been some works that studies RL with general function approximation. Russo and Roy (2013) and Osband and Roy (2014) initiated the study on the minimal structural assumptions that render sample-efficient learning by proposing a structural condition called Eluder dimension. Recently, several works have investigated different conditions for sample-efficient interactive learning, such as Bellman Eluder (BE) dimension (Jin et al., 2021), Bilinear classes (Du et al., 2021), Decision-Estimation Coefficient (DEC) (Foster et al., 2021), and Admissible Bellman Characterization (ABC) (Chen et al., 2023). A notable difference is that they impose assumptions on the value functions while we study function approximation on the transitions to ensure valid probability distributions. Moreover, the goal of these works is to study the conditions for sample-efficient reinforcement learning, but not focus on the computational efficiency.
@@
Multinomial Logit Bandits.  There are two types of multinomial logit bandits studied in the literature: the single-parameter model, where the parameter is a vector (Cheung and Simchi-Levi, 2017) and multiple-parameter model, where the parameter is a matrix (Amani and Thrampoulidis, 2021). We focus on the single-parameter model, which are more relevant to our setting. The pioneering work by Cheung and Simchi-Levi (2017) achieved a Bayesian regret of $\widetilde{\mathcal{O}}(\kappa^{-1}d\sqrt{T})$, where $T$ denotes the number of rounds in bandits. This result was further enhanced by subsequent studies (Oh and Iyengar, 2019, 2021; Agrawal et al., 2023). In particular, Périvier and Goyal (2022) significantly improved the dependence on $\kappa$, obtaining a regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa T}+\kappa^{-1})$ in the uniform revenue setting. Most prior methods required storing all historical data and faced computational challenge. To address this issue, the most recent work by Lee and Oh (2024) proposed an algorithm with constant computational and storage costs building on recent advances in multiple-parameter model (Zhang and Sugiyama, 2023). Their algorithm achieves the optimal regret of $\widetilde{\mathcal{O}}(d\sqrt{\kappa T}+\kappa^{-1})$ and $\widetilde{\mathcal{O}}(d\sqrt{T}+\kappa^{-1})$ under uniform and non-uniform rewards respectively. However, although the underlying models of MNL bandits and MDPs share similarities, the challenges they present differ substantially, and techniques developed for MNL bandits cannot be directly applied to MNL MDPs. For example, in MNL bandits, the objective is to select a series of assortments with varying sizes that maximize the expected revenue, whereas in MNL MDPs, the goal is to choose one action at each stage to maximize the cumulative reward. Thus, it is necessary to design new algorithms tailored for MDPs to address these unique challenges.
@@ 3 Problem Setup
Inhomogeneous, Episodic MDPs.  An inhomogeneous, episodic MDP instance can be denoted by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the length of each episode, $\mathbb{P}_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$ is the transition kernel with $\mathbb{P}_{h}(s^{\prime}\>|\>s,a)$ is being the probability of transferring to state $s^{\prime}$ from state $s$ and taking action $a$ at stage $h$, $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ is the deterministic reward function. A policy $\pi=\{\pi_{h}\}_{h=1}^{H}$ is a collection of mapping $\pi_{h}$, where each $\pi_{h}:\mathcal{S}\to\Delta(\mathcal{A})$ is a function maps a state $s$ to distributions over $\mathcal{A}$ at stage $h$. For any policy $\pi$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$, we define the action-value function $Q_{h}^{\pi}$ and value function $V_{h}^{\pi}$ as follows:
@@
where the expectation of $Q_{h}^{\pi}$ is taken over the randomness of the transition $\mathbb{P}$ and policy $\pi$. The optimal value function $V_{h}^{*}$ and action-value function $Q_{h}^{*}$ given by $V_{h}^{*}(s)=\sup_{\pi}V_{h}^{\pi}(s)$ and $Q_{h}^{*}(s,a)=\sup_{\pi}Q_{h}^{\pi}(s,a)$. For any function $V:\mathcal{S}\to\mathbb{R}$, we define $[\mathbb{P}_{h}V](s,a)=\mathbb{E}_{s^{\prime}\sim\mathbb{P}_{h}(\cdot\mid s,a)}V(s^{\prime})$.
Learning Protocol.  In the online MDP setting, the learner interacts with the environment without the knowledge of the transition kernel $\{\mathbb{P}_{h}\}_{h=1}^{H}$. We assume the reward function $\{r_{h}\}_{h=1}^{H}$ is deterministic and known to the learner. The interaction proceeds in $K$ episodes. At the beginning of episode $k$, the learner chooses a policy $\pi_{k}=\{\pi_{k,h}\}_{h=1}^{H}$. At each stage $h\in[H]$, starting from the initial state $s_{k,1}$, the learner observes the state $s_{k,h}$, chooses an action $a_{k,h}$ sampled from $\pi_{k,h}(\cdot\>|\>s_{k,h})$, obtains reward $r_{h}(s_{k,h},a_{k,h})$ and transits to the next state $s_{k,h+1}\sim\mathbb{P}_{h}(\cdot\>|\>s_{k,h},a_{k,h})$ for $h\in[H]$. The episode ends when $s_{H+1}$ is reached. The goal of the learner is to minimize regret, defined as
@@
Multinomial Logit (MNL) Mixture MDPs.  Although significant advances have been achieved for MDPs with linear function approximation, Hwang and Oh (2023) show that there exists a set of features such that no linear transition model can induce a valid probability distribution, which limits the expressiveness of function approximation. To overcome this limitation, they propose a new class of MDPs with multinomial logit function approximation. However, their work focuses on the homogeneous setting, where the transitions remain the same across all stages (i.e., $\mathbb{P}_{1}=...=\mathbb{P}_{H}$). In this work, we address the more general inhomogeneous setting, allowing transitions to vary across different stages. We introduce the formal definition of inhomogeneous MNL mixture MDPs below.
@@
An MDP instance $\mathcal{M}=(\mathcal{S},\mathcal{A},H,\{\mathbb{P}_{h}\}_{h=1}^{H},\{r_{h}\}_{h=1}^{H})$ is called an inhomogeneous, episodic $B$-bounded MNL mixture MDP if there exist a known feature mapping $\phi(s^{\prime}\>|\>s,a):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$ with $\lVert\phi(s^{\prime}\>|\>s,a)\rVert_{2}\leq 1$ and unknown vectors $\{\theta_{h}^{*}\}_{h=1}^{H}\in\Theta$ with $\Theta=\{\theta\in\mathbb{R}^{d},\lVert\theta\rVert_{2}\leq B\}$, such that for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$ and $s^{\prime}\in\mathcal{S}_{h,s,a}$, it holds that
@@
This model is consistent with models like neural networks (LeCun et al., 2015), where the feature $\phi$ is obtained by omitting the final layer, and $\theta_{h}^{*}$ represents the weights of the last layer. A final softmax layer is then applied to ensure that the output forms a valid probability distribution.
For any $\theta\in\mathbb{R}^{d}$, we define the induced transition as $p_{s,a}^{s^{\prime}}(\theta)=\frac{\exp(\phi(s^{\prime}\>|\>s,a)^{\top}\theta)}{\sum_{\widetilde{s}\in\mathcal{S}_{s,a}}\exp(\phi(\widetilde{s}\>|\>s,a)^{\top}\theta)}$. We then introduce the following two key problem-dependent quantities $\kappa$ and $\kappa^{*}$ that measure the effective non-linearity over the entire parameter space and around the ground truth parameter respectively.